OpenStudy (anonymous):

find any points on the curve x^2+xy=2y-8 where the line tangent to the curve is perpednicular to 2x+4y-1001=0?

OpenStudy (anonymous):

first, find the equation of the line perpendicular to 2x+4y-1001=0

OpenStudy (anonymous):

ok then what do i do?

OpenStudy (anonymous):

find the derivative of y wrt x

OpenStudy (anonymous):

\[x^2+xy=2y-8 \] \[2x+xy'+y=2y'\] solve for \(y'\)

OpenStudy (anonymous):

\[2x+4y-1001=0\] has slope \(-\frac{1}{2}\) perpendicular line will have slope \(2\) set \(y'=2\) and solve

OpenStudy (anonymous):

Ohh ok then Thanks!